OpenStudy (anonymous):

lim x->0 (tanx)/(4x)

OpenStudy (anonymous):

\[\lim_{\Theta \rightarrow 0} \frac{ \tan \theta }{ 4\theta }\]

OpenStudy (anonymous):

So, I would start by turning tan(theta) into sin/cos. Then keep in mind that we have a common trig limit that is taught, namely: \[\lim_{x \rightarrow 0}\frac{ sinx }{ x }= 1\]So you want to manipulate what you have to try and match that common limit. Would you have an idea how to go about that?

OpenStudy (anonymous):

\[\lim_{\theta \rightarrow 0} \frac{ 1 }{ 4\cos \theta }\] Can I do that?

OpenStudy (anonymous):

If I can, then 1/4

OpenStudy (jdoe0001):

\(\bf \lim_{\theta \to 0} \cfrac{ \tan \theta }{ 4\theta } \\ \quad \\ \cfrac{ \tan \theta }{ 4\theta }\implies \cfrac{\frac{sin(\theta)}{cos(\theta)}}{4\theta}\implies \cfrac{sin(\theta)}{cos(\theta)}\cdot \cfrac{1}{4\theta} \\ \quad \\ \cfrac{sin(\theta)}{cos(\theta)4\theta}\implies \cfrac{sin(\theta)}{\theta}\cdot \cfrac{1}{4cos(\theta)}\)

OpenStudy (anonymous):

Perfect. Thanks.

OpenStudy (anonymous):

He did the work I was typing out, lol.

OpenStudy (anonymous):

But you're correct, 1/4 :P

OpenStudy (jdoe0001):

yw